Match the confidence level with the confidence interval for the population mean.

1. ?bar±2.575(?/√n)

2. ?bar±1.645(?/√n)

3. ?bar±1.282(?/√n)

A. 80% B. 90% C. 99%
Answer #1

solution:

1. ?bar±2.575(?/√n)

At 99% confidence level the z is,

= 1 - 99%

= 1 - 0.99 = 0.01

/2 = 0.005

Z/2 = 2.576

2. ?bar±1.645(?/√n)

At 90% confidence level

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645

3.

?bar±1.282(?/√n)

At 80% confidence level

= 1 - 80%  

= 1 - 0.80 =0.20

/2 = 0.10

Z/2 = Z0.10 = 1.282